find rational number r=m/n . Such that √5 < r < √6


1
Expert's answer
2021-01-12T17:16:26-0500

The distance between the numbers "\sqrt{5}" and "\sqrt{6}" is equal to

"\sqrt{6}-\sqrt{5} = \frac{(\sqrt{6}-\sqrt{5})(\sqrt{6}+\sqrt{5})}{\sqrt{6}+\sqrt{5}} = \frac{1}{\sqrt{6}+\sqrt{5}}>\frac15"

Therefore, if "n\geq 5" , then the length of the interval ("\sqrt{5}n,\sqrt{6}n" ) is equal to "\sqrt{6}n-\sqrt{5}n >1" and this interval must contain an integer m.

We have "\sqrt{5}n < m <\sqrt{6}n" and, therefore, the ratio m/n belongs to the interval "(\sqrt{5},\sqrt{6})".

To find the ratio m/n in the interval "(\sqrt{5},\sqrt{6})" we should seek a full-squared integer m2 in the interval (5n2, 6n2).

Let's take n=5.

"5n^2 = 5\cdot 5^2 = 125"

"6n^2 = 6\cdot 5^2 = 150"

The only full-squared integer between 125 and 150 is 144=122=m2.

Therefore, r=m/n = 12/5 = 2.4 satisfies the original condition.